# ARFlow: Human Action-Reaction Flow Matching with Physical Guidance

### v2

@@ 7 Parameter analysis of guidance strength
Table 7.1: Parameter analysis of guidance strength on the online, unconstrained setting on NTU120-AS. Bold indicates the best result.
@@
Real 0.09 38.38% 3.33
     $\lambda_{\text{pene}}$ 0 7.89 17.44% 1.55
1 8.29 16.01% 1.43
2 8.51 13.30% 1.15
5 8.65 9.60% 0.81
10 8.75 8.56% 0.76
15 9.04 8.50% 0.75
8 Details of our framework
@@
Figure 8.1: Human perceptual study results on NTU120-AS.
Figure 8.2: Human perceptual study results on Chi3D-AS.
9 Implementation details
Our ARFlow model is trained with $T=1,000$ timesteps using a classifier-free approach [23]. The number of decoder layers is 8 and the latent dimension of the Transformer tokens is 512. The batch size is configured as 32 for NTU120-AS and 16 for Chi3D-AS. The interaction loss weight is set to $\lambda_{\text{inter}}=1$. Each model is trained for 500K steps on single NVIDIA 4090 GPU within 48 hours. During inference, unless otherwise stated, we employ 5-timestep sampling for all the diffusion-based and our models in our experiments as [69] for a fair comparison. For the physical constraint guidance, we set the safe distance $\zeta=1$ and $\lambda_{\text{pene}}=10$
10 Details of the metric calculations
@@
Figure 10.1: Visualization comparison of the effects before and after using physical constraint guidance. Blue for actors and Green for reactors.
11 Extra qualitative results
We show the generated motions of our method against others in Fig. 11.2. We highlight the implausible motions in rectangle marks, it is clear that our method learns the correct reactions and avoids human-human inter-penetrations as much as possible.
Figure 11.1: Failure case of our method.
Figure 11.2: The extra qualitative experiment. Blue for actors and Green for reactors.
Failure case. We also show the failure cases of our motion generation pipeline in Fig. 11.1. Our model cannot guarantee absolute physical authenticity, for example, ensuring that the hand contacts but does not penetrate during handshaking. Incorporating more sophisticated physical constraints may solve the failure cases and be considered in future.
12 User Study
@@
The results are shown in Fig. 8.1 and Fig. 8.2. Our results were preferred over the other state-of-the-art and are even competitive with ground truth motions.

### v3

@@ Appendix C Influence of sampling randomness
Table C.1: Randomness Influence studies on the online, unconstrained setting on the NTU120-AS dataset. Bold indicates the best result in our method.
Appendix D Parameter analysis of guidance strength and weight factor
We conduct a parameter analysis of guidance strength in Tab. D.1. The result of the experiments show that as the guiding strength increases, the degree of penetration between actors and reactors decreases significantly, while FID increases slightly. This is because the ground truth itself has a certain degree of penetration. Thus, this task requires our new metrics and FID to collaborate in evaluating the quality of the generated results. When the guidance strength increases to a certain extent, the decrease in penetration degree is no longer significant. Therefore, we ultimately choose $\lambda_{\text{pene}}=2$. Our method achieves the lowest penetration level while maintaining the best generation quality. As for the weight factor, the results show that the minimum value of FID does not occur at the endpoints, thus demonstrating the effectiveness of our weighting method.
Table D.1: Parameter analysis of guidance strength and weight factor on the online, unconstrained setting on NTU120-AS. Bold indicates the best result.
@@ Appendix D Parameter analysis of guidance strength and weight factor
Real 0.09 21.96% 5.35
     $\lambda_{\text{pene}}$ 0 7.89 8.39% 3.26
1 7.98 5.80% 1.15
2 8.20 3.54% 0.68
5 8.49 1.22% 0.13
10 9.41 0.78% 0.21
     $w$ 0 8.37 2.71% 0.35
0.1 8.30 2.78% 0.36
0.3 8.19 2.96% 0.37
0.5 8.11 3.12% 0.41
0.7 8.07 3.23% 0.53
0.9 8.08 3.32% 0.64
1 8.20 3.54% 0.68
Appendix E Details of our framework
@@
Appendix F Implementation details
Our ARFlow model is trained with $T=1,000$ timesteps using a classifier-free approach [25]. The number of decoder layers is 8 and the latent dimension of the Transformer tokens is 512. The batch size is configured as 32 for NTU120-AS, InterHuman-AS and 16 for Chi3D-AS. The interaction loss weight is set to $\lambda_{\text{inter}}=1$. Each model is trained for 500K steps on single NVIDIA 4090 GPU within 48 hours. During inference, unless otherwise stated, we employ 5-timestep sampling for all the diffusion-based and our models in our experiments as [72] for a fair comparison. For the physical constraint guidance, we set the safe distance $\zeta=0.5$, $\lambda_{\text{pene}}=2$ and $w=0.7$
Figure F.1: Human perceptual study results on NTU120-AS.
Figure F.2: Human perceptual study results on Chi3D-AS.
Appendix G Details of the metric calculations
@@
Appendix H User Study
@@
The results are shown in Fig. F.1 and Fig. F.2. Our results were preferred over the other state-of-the-art and are even competitive with ground truth motions.
Figure H.1: Visualization comparison of the effects before and after using physical constraint guidance. Blue for actors and Green for reactors.
Appendix I Extra qualitative results
We show the generated motions of our method against others in Fig. I.1. We highlight the implausible motions in rectangle marks, it is clear that our method learns the correct reactions and avoids human-human inter-penetrations as much as possible.
Figure I.1: The extra qualitative experiment. Blue for actors and Green for reactors.
Figure I.2: Failure case of our method.
Failure case. We also show the failure cases of our motion generation pipeline in Fig. I.2. Our model cannot guarantee absolute physical authenticity, for example, ensuring that the hand contacts but does not penetrate during handshaking. Incorporating more sophisticated physical constraints may solve the failure cases and be considered in future.
Appendix J Broader Impacts
Our model demonstrates significant potential for AR/VR and gaming applications by enabling the generation of plausible human reactions. Beyond virtual environments, the proposed approach provides an innovative technical pathway for real-world human-robot interaction, where motion patterns can be transferred to robotic systems through motion remapping technology. Although this advancement may inspire future research, we acknowledge potential misuse risks similar to other generative models, warranting ethical considerations as the technology develops.